|
| ||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||||
| Stack Data | Did not stack | ||||||||||||
|---|---|---|---|---|---|---|---|---|---|---|---|---|---|
| Problem Reports | ||||
|---|---|---|---|---|
| PR Number | Times | Time Lost | Description | Response |
| 3992 | 01:32-04:17 | 02:45 | Could not get an image from the MARS camera. Jim, George, and Mike swapped cables and fibers in the tower, control room, and UER until we got an abnormal signal on the AOA camera. Suspected camera failure. | |
| 3993 | 03:00-04:17 | 01:17 | HEFI image is scrambled and cuts out and in accompanied by incessant chirps in the control room. HEFI is unusable in this state. | |
| Ops Comments | none |
|---|
| TO Comments | The night was mainly devoted to figuring out problems with HEFI/MARS video. No stacking, testing. |
|---|
| RA Comments | none |
|---|
| Observations | |||||||||||||||
|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|
| File | Target/QID | Exp Start | Exposure | Overhead | Program | Instrument Configuration | Data Type | IQ | Dimm | SkySBrt | Skytrans | SNR | Comments | Charge | |